OpenStudy (anonymous):

create your own two step equation with all real numbers as the solution. show the steps to solving the equation to prove the solution all real numbers. help?

OpenStudy (ashleyisakitty):

Tell me what you think the first steps to this would be

OpenStudy (mww):

All real numbers as a solution implies that no matter what value of x you use, LHS of the equation = RHS of the equation. The only way this would work is if the equations were identical. For such equations, when you solve them, you eventually end up with a 0 = 0 result. So construct two equations which are identical, or equivalent and make them equal to each other.

OpenStudy (anonymous):

ok! thxs!

OpenStudy (anonymous):

so like 2x+3=2x+3?

OpenStudy (mww):

But try not to make them too obvious...you want to write two equations, but one of them is an alternative way to write the equation. eg. x^2 - 9 = (x +3)(x -3)

OpenStudy (mww):

that is still correct though

OpenStudy (anonymous):

oh ok! thanks

OpenStudy (mww):

The idea is that you can see no matter what x you use, the LHS will always be the RHS because the expressions are identical. So graphically, if you were to draw them, you expect them to pass each other infinite times (same graph)